Solve the logarithmic equation.reject any value of x that is not in the domain of the original logarithmic expression. log (x + 16) = log x + log 16
Answer by josgarithmetic(39617)

Righthand side is log(x*16), so the equation shown can be same as log(x+16)=log(16*x). Meaning is x+16=16x.
x-16x=-16
-15x=-16
x=16/15